OpenStudy (anonymous):

Solve: \[(6-\sqrt{7})(5+\sqrt{7})\]

OpenStudy (anonymous):

you gotta multiply this out: \[30+6\sqrt{7}-5\sqrt{7}-7\], then simplify

OpenStudy (anonymous):

23+1√7?

OpenStudy (anonymous):

sure

jimthompson5910 (jim_thompson5910):

\[\large (6-\sqrt{7})(5+\sqrt{7})\] \[\large 6*5+6*\sqrt{7}+(-\sqrt{7})(5)+(-\sqrt{7})(\sqrt{7})\] \[\large 30+6\sqrt{7}-5\sqrt{7}-\sqrt{7*7}\] \[\large 30+6\sqrt{7}-5\sqrt{7}-\sqrt{49}\] \[\large 30+6\sqrt{7}-5\sqrt{7}-7\] \[\large 23+\sqrt{7}\] So \[\large (6-\sqrt{7})(5+\sqrt{7})=23+\sqrt{7}\]

OpenStudy (anonymous):

Thank you for explaining Jim :)